OpenStudy (lin.ivory):

Simplify: (((a-4)^4)-((b+4)^4))/(((a-4)^3)-((b+4)^3))

OpenStudy (anonymous):

\[((a-4)^4)-((b+4)^4))/(((a-4)^3)-((b+4)^3)) \] do you mean this?

OpenStudy (lin.ivory):

yes!!

OpenStudy (anonymous):

\[\frac{(a-4)^4)-((b+4)^4)}{ ((a-4)^3)-((b+4)^3) }\] \[\frac{((a-4)^2)^2-((b+4)^2)^2}{ ((a-4)^3)-((b+4)^3) }\] and use the formulaes x^2 - y^2 = (x + y)( x - y ) and x^3 - y^3=(x-y)(x^2+xy+y^2)

OpenStudy (lin.ivory):

but why does \[X ^{2} - Y ^{2} = (x-y)(x+y)\]?

OpenStudy (lin.ivory):

OHHHHHHHHHHHH nevermind!!

OpenStudy (lin.ivory):

AHAHAHHAHAh kkkz!!! THANKS!